OpenStudy (anonymous):

Help me find the area and perimeter of a rhombus with the side length of 5.385

OpenStudy (anonymous):

I either need a pic, or a measure of the four angles of the rhombus.

OpenStudy (anonymous):

P=4(5.385)

OpenStudy (anonymous):

This is a picture of my question, but I have already found the length of each side. It is 5.385 @Zelda

OpenStudy (anonymous):

@Jonask I have the p, all I need is the area. Idk how though?

OpenStudy (anonymous):

this picture is talking about parrallelogram

OpenStudy (anonymous):

I know but the shape that was drawn was a rhombus.

OpenStudy (anonymous):

look work it out & you will see.

OpenStudy (anonymous):

you are talking about a rhombus???

OpenStudy (anonymous):

The area of a rhombus is 1/2 * the product of the length of the diagonals.

OpenStudy (anonymous):

Yes. & so how do I find the diagonal?

OpenStudy (anonymous):

distance formular\[D=\sqrt{x_1-x_2)^2+(y_1-y_2)^2}\]

OpenStudy (anonymous):

(-1,2) and (2,-1)

OpenStudy (anonymous):

And the other diagonal is (4,4) and (-3,3)

OpenStudy (anonymous):

But don't I want the diagonal ?

OpenStudy (anonymous):

Yes you do. You need that for area.

OpenStudy (anonymous):

So do A = .5 * D1 * D2

OpenStudy (anonymous):

Im still confused. To find the area you need to do 1/2 d1 x d2?

OpenStudy (anonymous):

But idk how to find diagonal 1?

OpenStudy (anonymous):

Yes, so you need both diagonals lengths.

OpenStudy (anonymous):

One sec.

OpenStudy (anonymous):

\[A=\frac{d_1d_2}{2}\] \[d_1=3\sqrt{2} \space d_2=5\sqrt{2}\]

OpenStudy (anonymous):

So, I attached a Pic, those two are your diagonals, so use the distance formula to get the Diagonal length for (4,4) to (-3,-3) D1, and then (-1,2) to (2,-1) D2. Once you have the lengths: Use this to find Area: \[A = \frac{ D1 * D2 }{ 2 }\]

OpenStudy (anonymous):

okay D1 = 9.899 & D2 = 4.242 for the lengths? @Zelda

OpenStudy (anonymous):

Area = 20.995?

OpenStudy (anonymous):

@Jonask is my area right?

OpenStudy (anonymous):

D1 = 4,4 & 3,3 you need to find the distance of that right?

OpenStudy (anonymous):

sry i was using wrong,i used -3,3 instead of -3,-3 cordinates you are correct

OpenStudy (anonymous):

So my area is 21.54

OpenStudy (anonymous):

but to be accurate use the roots\[\frac{7\sqrt{2}(3\sqrt{2})}{2}=21\]

OpenStudy (anonymous):

Alright thank you!

OpenStudy (anonymous):

No problem, your welcome.